# Can AI writing be salvaged? Mitigating Idiosyncrasies and Improving Human-AI Alignment in the Writing Process through Edits

### v3

@@ Abstract.
LLM-based applications are helping people write, and LLM-generated text is making its way into social media, journalism, and our classrooms. However, the differences between LLM-generated and human-written text remain unclear. To explore this, we hired professional writers to edit paragraphs in several creative domains. We first found these writers agree on undesirable idiosyncrasies in LLM-generated text, formalizing it into a seven-category taxonomy (e.g. clichés, unnecessary exposition). Second, we curated the LAMP corpus: 1,057 LLM-generated paragraphs edited by professional writers according to our taxonomy. Analysis of LAMP reveals that none of the LLMs used in our study (GPT4o, Claude-3.5-Sonnet, Llama-3.1-70b) outperform each other in terms of writing quality, revealing common limitations across model families. Third, we explored automatic editing methods to improve LLM-generated text. A large-scale preference annotation confirms that although experts largely prefer text edited by other experts, automatic editing methods show promise in improving alignment between LLM-generated and human-written text.
@@ 1. Introduction
While LLM-based writing assistants have the potential to improve writing quality and increase author productivity, they also introduce an algorithmic monoculture (Kleinberg and Raghavan, 2021). Padmakumar and He (2023a) and Anderson et al. (2024a) discuss how writing with LLMs unintentionally reduces content diversity, leading to homogenization. This homogenization occurs not just at the semantic level but also at syntactic (structural), lexical (Kobak et al., 2024), and stylistic levels (Shaib et al., 2024). For instance, prior work from Ippolito et al. (2022b); Chakrabarty et al. (2024b) has shown how LLM-generated text is often hackneyed and rife with clichés, while failing to demonstrate rhetorical complexity and often revealing the subtext —a phenomenon known as “telling instead of showing” (Chakrabarty et al., 2024a). Additionally, LLM-generated texts are typically full of redundant exposition, overwrought metaphors, and florid descriptions due to verbosity bias during preference labeling (Saito et al., 2023).
@@
Current AI-assisted writing tools are powered by pre-trained language models that are refined through Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019). RLHF transforms human preferences into training data to guide language models toward desired outcomes. The most common type of feedback used with RLHF is binary preferences between pairs of examples sampled from one or more language models (Christiano et al., 2017) (See Fig 1 Traditional Alignment). However, this approach has a drawback. The paired outputs may differ in numerous ways and could be equally flawed in containing idiosyncrasies. Asking an annotator to choose between two undesirable outputs does not improve alignment [^1] In the current design of RLHF annotators are not allowed to not pick either. (Casper et al., 2023; Ghosh et al., 2024). We argue that alignment training needs to be aware of how desirable any individual response is, regardless of its preference relationship. Editing undesirable portions of a response can be seen as an effective mechanism for enhancing alignment (See Fig 1 Alignment via Edits). An LLM-generated response that has been edited typically contains fewer undesirable traits and can be paired with the original LLM-generated response for preference ranking ($edited>original$). The challenge however lies in consistently identifying and implementing edits that enhance quality while aligning with human preferences for effective writing. Prompting techniques that encourage the model to self-edit (Madaan et al., 2024) have shown promise however they do not work for long form or paragraph level writing (Pan et al., 2024). The primary reason for this can be accounted to the fact that LLMs don’t inherently know what aspects of the writing needs improvement, how extensively they should edit, or how to make changes that align with any given writer’s expertise level and style.
@@
To address these problems, we first create a comprehensive taxonomy of edit categories based on expert writing practices. We then recruit 18 writers to edit LLM-generated text using categories from our taxonomy. We define edits as changes that alter, replace, or refine specific phrases, clauses, or sentences within a larger text. We restrict our focus to generating text in literary fiction and creative non-fiction, as these genres challenge LLMs with their creativity, emotional nuance, and sophisticated language use. We focus on paragraph-level edits as they balance granularity and scope, reducing costs and annotator fatigue. Paragraphs capture style and context better than sentences, enabling more cohesive improvements. Given LLMs’ limitations in long-term discourse coherence, paragraph-level enhancements facilitate human-AI collaboration. This approach allows humans to guide the overall structure and flow while AI handles lower-level details (Dhillon et al., 2024). Finally, while expert writer’s edits are valuable for identifying idiosyncrasies, this approach may not scale for large text volumes. To address this, we design few-shot prompts that use writer’s edits to identify problematic spans in LLM-generated paragraphs and suggest improvements. This method aims to enhance overall paragraph quality at scale. To summarize our contributions:
@@
  • We conduct an empirical investigation that tests if LLMs can automatically detect and rewrite their own idiosyncrasies. Our statistically significant results show an encouraging preference trend $Writer-edited>LLM-edited>LLM-generated$ suggesting that edits improve human-AI alignment in the writing process.
@@
Our code, data and experimental setup can be found at  [^2] https://github.com/salesforce/creativity_eval/.
@@ 2.1. Text Editing in HCI
Text editing is the process of modifying written content using specialized software. HCI research on text editing aims to improve digital writing tools’ efficiency and usability. Word processors have long allowed flexible editing functions (Pea and Kurland, 1987). Systems like Soylent (Bernstein et al., 2015), MicroWriter (Teevan et al., 2016), and WearWrite (Nebeling et al., 2016) developed interfaces for crowd-based editing, focusing on task breakdown, cost management, and minimizing delays. Robertson and Black (1983) proposes a goal-fate analysis model for text editing behavior, supported by data showing distinct plan units in editing tasks, with potential for intention-based user assistance. Tyler et al. (1982) investigated text editing skill progression and effective training methods. Rosson (1983) examined the impact of experience on editing behavior, questioning if users naturally develop optimal strategies or plateau, noting that experienced users tend to develop more efficient editing heuristics than novices. Reza et al. (2024) present ABScribe, a novel interface that streamlines the process of generating and comparing writing variations using Large Language Models, addressing challenges in existing text-editing workflows and improving writer’s efficiency and satisfaction. Zhou and Sterman (2024) suggest that imperfect AI text suggestions can promote deeper engagement in rewriting, potentially preserving the writer’s authenticity and creative ownership. Park and Lee (2023) found that providing rationales for edits in collaborative writing was generally preferred by participants, despite no significant differences in survey results. This led to design recommendations for effective collaboration. Dang et al. (2022) propose a text editor with continuously updated paragraph-wise summaries as margin annotations to help users plan, structure, and reflect on their writing process.Laban et al. (2023a) introduce InkSync, an LLM-based editing interface suggesting executable document edits. It uses a three-stage approach (Warn, Verify, Audit) to reduce factual errors and enhance editing accuracy, efficiency, and user experience compared to standard chat interfaces. In contrast to existing research, we focus on edits as a method to improve human-AI alignment in writing assistance. Our work characterizes the undesirable aspects of AI writing informed by expert consensus and designs an approach to mitigate these through text editing.
@@ 2.4. Human AI alignment in Writing
Lee et al. (2024) highlight how AI tools have transformed writing processes, establishing new criteria for future AI writing assistants. In a similar vein Li et al. (2024) reveal that while users benefit from AI assistance in productivity and confidence, potential drawbacks include decreased accountability and diversity in writing. LLMs used in writing assistance can significantly influence human-authored content. Hohenstein and Jung (2020) found LLM-generated text suggestions can affect human writer’s emotional tone. Arnold et al. (2020) showed predictive text encourages predictable writing. Anderson et al. (2024b) and Laban et al. (2023a) found LLMs like ChatGPT helped users generate more detailed ideas, but outputs were less semantically distinct across users (Padmakumar and He, 2023b), and participants felt less responsible for their produced ideas. Recent work from Pan et al. (2024) demonstrates language models can enhance outputs via feedback. However, methods like Iterative Self-refinement scenarios, using another language model as an evaluator, may result in reward hacking, where the model exploits the evaluator’s flaws. For alignment training, it’s crucial to consider the absolute desirability of each potential response, not just how responses compare to one another in terms of preference. Towards this in our work, we create pairs consisting of an initial LLM-generated response and its refined counterparts that by nature is more contrastive (or closely comparable). Our results show that such a pairing results in improved alignment and agreement during preference ranking.
@@ 3. DESIGN CONSIDERATIONS TO IMPROVE AI WRITING
The secret to good writing is good editing (to Edit Your Own Writing, 2024). It’s what separates hastily written, randomly punctuated, incoherent rants from learned polemics and op-eds, and cringe-worthy fan fiction from a critically acclaimed novel (to Edit Your Own Writing, 2024). In this section we outline the design principles and desiderata that guided our approach in improving AI writing through textual edits.
@@
This principle emphasizes creating a comprehensive taxonomy of edit categories (Faigley and Witte, 1981) rooted in expert editor and writer’s practices. Prior work has shown that experts and novices define revising in very different ways with experts attending more systematically to different aspects of the text than novices (Bracewell et al., 1978; Sommers, 1980; Tyler et al., 1982; Rosson, 1983). By developing such a taxonomy, we aim to provide an approach to analyzing and enhancing LLM-generated text. It also allows for a more granular understanding of the specific areas where AI writing may fall short and enables targeted improvements. Sommers (1980) found that “experienced writers have a second objective; a concern for their readership”. Grounding the taxonomy in expert writing practices ensures that the edits align with the standards of high-quality writing and is acceptable to its readers. Finally this principle also acknowledges the complexity of the editing process, recognizing that different categories of edits may be required at various levels of the text, from sentence-level corrections to broader structural changes (Hayes et al., 1987; Sommers, 1980; Yang et al., 2017).
@@
Preserving core meaning and intent of the original text is crucial to maintain coherence and faithfulness to the initial ideas. On the other hand, introducing substantive semantic changes is often required to adhere to the quality and characteristics of good writing. Prior work on edit taxonomies focuses on low-level syntactic operations (Faigley and Witte, 1981) or semantic edits (Yang et al., 2017; Daxenberger and Gurevych, 2013; Laban et al., 2023b) tailored to specific websites like Wikipedia. LLM-generated text often benefits from syntactic edits. These edits (primarily meaning preserving) enhance readability by diversifying sentence structures, expanding vocabulary choices, and minimizing repetitive phrasing. Consequently semantic edits (both meaning preserving and changing) in AI writing are important for enhancing specificity or reducing unnecessary flourishes and clichés that can otherwise obscure meaning. Our methodology aims to navigate the tension between maintaining original meaning and introducing necessary improvements to mitigate AI-specific writing quirks.
@@
Current AI writing systems are developed using pre-trained language models (LMs) refined through human interaction, employing supervised learning and reinforcement learning (RL) techniques. Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019) is a key approach, transforming human input into training data to guide LMs toward desired outcomes. The most common type of feedback used with RLHF is binary preferences between pairs of examples sampled from one or more Language Models (Christiano et al., 2017). However, a learned preference ordering can fail to converge to the true one when the desirability of examples depends on noise (Gao et al., 2024). To tackle this we propose edits as a mechanism for enhancing alignment. An LLM-generated response that has been edited typically contains fewer undesirable traits and can be paired with the original LLM-generated response for preference ranking ($edited>original$). This method of incorporating edits for improved preference data collection has also been adopted by contemporaneous work from Meta AI (Dubey et al., 2024) for training Llama3.1 as well as Contextual AI (D’Oosterlinck et al., 2024).
@@ 4.1. Collecting Instruction and Response Pairs
Finally, we use the generated instructions to elicit responses from three state-of-the-art LLMs: OpenAI’s GPT-4o (OpenAI, 2024), Anthropic’s Claude-3.5-Sonnet (Anthropic, 2024), and Meta’s Llama 3.1-70b (Dubey et al., 2024). Each LLM is used to generate responses to one-third of our instruction data. We ensure that each LLM responds to instruction across all domains in equal proportion. To generate high-quality responses, we provide each LLM with the writing instruction, as well as the genre and source, and instruct it to adhere to the style of the venue. The prompt further specifies: “Try your best to be original, avoiding clichés or overused tropes. Do not use ornamental language and focus on nuance, simplicity, and subtext” (See Prompt in Appendix A Table 13). Through this process we obtain 1,057 writing $<$instructions, response$>$ pairs, with responses averaging 205 words. This collection of instructions and LLM-generated responses serves as the foundation for the three studies we conducted: the formative study, the full-scale editing annotation, and the preference annotation.
@@ 4.4.6. Awkward Word Choice and Phrasing
Awkward phrasing can significantly reduce writing quality, often confusing or disengaging readers. This issue typically involves misused or disproportionate use of certain words (Kobak et al., 2024), unclear pronoun references, or an overuse of passive voice. In an email W1 pointed out “Another little observation to share: a very common phrasing in these excerpts is ‘seem to _(verb)_’. This is not technically wrong it’s just inelegant, something many writing teachers have told me to avoid. Unless there is some specific uncertainty or doubt about the verb action, it’s always preferred to just use the verb without ‘seem’ (ex. from the current excerpt I have up: ‘amplified’ is better than ‘seemed to amplify’)”. Editing plays a crucial role in refining these elements. Through careful revision, writers can identify and replace imprecise or ill-fitting words with more appropriate alternatives, ensuring each term accurately conveys the intended meaning.

### v5

@@ Abstract.
LLM-based applications are helping people write, and LLM-generated text is making its way into social media, journalism, and our classrooms. However, the differences between LLM-generated and human-written text remain unclear. To explore this, we hired professional writers to edit paragraphs in several creative domains. We first found these writers agree on undesirable idiosyncrasies in LLM-generated text, formalizing it into a seven-category taxonomy (e.g. clichés, unnecessary exposition). Second, we curated the LAMP corpus: 1,057 LLM-generated paragraphs edited by professional writers according to our taxonomy. Analysis of LAMP reveals that none of the LLMs used in our study (GPT4o, Claude-3.5-Sonnet, Llama-3.1-70b) outperform each other in terms of writing quality, revealing common limitations across model families. Third, building on existing work in automatic editing we evaluated methods to improve LLM-generated text. A large-scale preference annotation confirms that although experts largely prefer text edited by other experts, automatic editing methods show promise in improving alignment between LLM-generated and human-written text.
@@ 1. Introduction
While LLM-based writing assistants have the potential to improve writing quality and increase author productivity, they also introduce an algorithmic monoculture (Kleinberg and Raghavan, 2021). Padmakumar and He (2023) and Anderson et al. (2024a) discuss how writing with LLMs unintentionally reduces content diversity, leading to homogenization. This homogenization occurs not just at the semantic level but also at syntactic (structural), lexical (Kobak et al., 2024), and stylistic levels (Shaib et al., 2024). For instance, prior work from Ippolito et al. (2022b); Chakrabarty et al. (2024b) has shown how LLM-generated text is often hackneyed and rife with clichés while failing to demonstrate rhetorical complexity and often revealing the subtext —a phenomenon known as “telling instead of showing” (Chakrabarty et al., 2024a). Additionally, LLM-generated texts are typically full of redundant exposition, overwrought metaphors, and florid descriptions due to verbosity bias during preference labeling (Saito et al., 2023).
@@
Current AI-assisted writing tools are powered by pre-trained language models that are refined through reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019). RLHF transforms human preferences into training data to guide language models toward desired outcomes. The most common type of feedback used with RLHF is binary preferences between pairs of examples sampled from one or more language models (Christiano et al., 2017) (See Fig 1 Traditional Alignment). However, this approach has a drawback. The paired outputs may differ in numerous ways and could be equally flawed in containing idiosyncrasies. Asking an annotator to choose between two undesirable outputs does not improve alignment [^1] In the current design of RLHF annotators are not allowed to not pick either. (Casper et al., 2023; Ghosh et al., 2024). We argue that alignment training needs to be aware of how desirable any individual response is, regardless of its preference relationship. Editing undesirable portions of a response can be seen as an effective mechanism for enhancing alignment (See Fig 1 Alignment via Edits). An LLM-generated response that has been edited typically contains fewer undesirable traits and can be paired with the original LLM-generated response for preference ranking ($edited>original$). The challenge, however, lies in consistently identifying and implementing edits that enhance quality while aligning with human preferences for effective writing. Prompting techniques that encourage the model to self-edit (Madaan et al., 2024) have shown promise; however, they do not work for long-form or paragraph-level writing (Pan et al., 2024a). The primary reason for this is that LLMs do not inherently know what aspects of the writing need improvement, how extensively they should edit, or how to make changes that align with any given writer’s expertise level and style.
@@
To address these problems, we first create a comprehensive taxonomy of edit categories based on expert writing practices. We then recruit 18 writers to edit LLM-generated text using categories from our taxonomy. We define edits as changes that alter, replace, or refine specific phrases, clauses, or sentences within a larger text. We restrict our focus to generating text in literary fiction and creative non-fiction, as these genres challenge LLMs with their creativity, emotional nuance, and sophisticated language use. We focus on paragraph-level edits, as they balance granularity and scope, reducing costs and annotator fatigue. Paragraphs capture style and context better than sentences, enabling more cohesive improvements. Given LLMs’ limitations in long-term discourse coherence, paragraph-level enhancements facilitate human-AI collaboration. This approach allows humans to guide the overall structure and flow, while AI handles the lower-level details (Dhillon et al., 2024). Finally, while expert writer’s edits are valuable for identifying idiosyncrasies, this approach may not scale for large text volumes. To address this, we design few-shot prompts that use writer’s edits to identify problematic spans in LLM-generated paragraphs and suggest improvements. This method aims to enhance overall paragraph quality at scale. Our work adds to the existing rich literature (Liu et al., 2022; Jiang et al., 2022; Gao et al., 2024b; Yao et al., 2023) on using completion edits to improve alignment in different tasks. To summarize our contributions:
@@
  • Building on prior work Hayes et al. (1987); Scardamalia (1983) we conduct an empirical investigation that tests if LLMs can automatically detect and rewrite their own idiosyncrasies. Our statistically significant results show an encouraging preference trend $Writer-edited>LLM-edited>LLM-generated$ suggesting that edits improve human-AI alignment in the writing process.
@@
Our code, data, and experimental setup is available at  [^2] https://github.com/salesforce/creativity_eval/tree/main/Writing_Alignment.
@@ 2.1. Text Editing in HCI
Text editing is the process of modifying written content using specialized software. HCI research on text editing aims to improve digital writing tools’ efficiency and usability. Word processors have long allowed flexible editing functions (Pea and Kurland, 1987). Systems like Soylent (Bernstein et al., 2015), MicroWriter (Teevan et al., 2016), and WearWrite (Nebeling et al., 2016) developed interfaces for crowd-based editing, focusing on task breakdown, cost management, and minimizing delays. Robertson and Black (1983) proposes a goal-fate analysis model for text editing behavior, supported by data showing distinct plan units in editing tasks, with the potential for intention-based user assistance. Tyler et al. (1982) investigated text editing skill progression and effective training methods. Rosson (1983) examined the impact of experience on editing behavior, questioning if users naturally develop optimal strategies or plateau, noting that experienced users tend to develop more efficient editing heuristics than novices. Reza et al. (2024) present ABScribe, a novel interface that streamlines the process of generating and comparing writing variations using Large Language Models, addressing challenges in existing text-editing workflows and improving writer’s efficiency and satisfaction. Zhou and Sterman (2024) suggest that imperfect AI text suggestions can promote deeper engagement in rewriting, potentially preserving the writer’s authenticity and creative ownership. Park and Lee (2023) found that providing rationales for edits in collaborative writing was generally preferred by participants, despite no significant differences in survey results. This led to design recommendations for effective collaboration. Dang et al. (2022) propose a text editor with continuously updated paragraph-wise summaries as margin annotations to help users plan, structure, and reflect on their writing process.Laban et al. (2023a) introduce InkSync, an LLM-based editing interface suggesting executable document edits. It uses a three-stage approach (Warn, Verify, Audit) to reduce factual errors and enhance editing accuracy, efficiency, and user experience compared to standard chat interfaces. In contrast to existing research, we focus on edits as a method to improve human-AI alignment in writing assistance. Our work characterizes the undesirable aspects of AI writing informed by expert consensus and designs an approach to mitigate these through text editing.
@@ 2.4. Human AI alignment in Writing
Lee et al. (2024) highlight how AI tools have transformed writing processes, establishing new criteria for future AI writing assistants. In a similar vein Li et al. (2024) reveal that while users benefit from AI assistance in productivity and confidence, potential drawbacks include decreased accountability and diversity in writing. LLMs used in writing assistance can significantly influence human-authored content. Hohenstein and Jung (2020) found LLM-generated text suggestions can affect a human writer’s emotional tone. Arnold et al. (2020) showed predictive text encourages predictable writing. Anderson et al. (2024b) and Laban et al. (2023a) found LLMs like ChatGPT helped users generate more detailed ideas, but outputs were less semantically distinct across users (Padmakumar and He, 2023), and participants felt less responsible for their produced ideas. Recent work from Pan et al. (2024a) demonstrates language models can enhance outputs via feedback. However, methods like Iterative Self-refinement scenarios, using another language model as an evaluator, may result in reward hacking, where the model exploits the evaluator’s flaws. For alignment training, it’s crucial to consider the absolute desirability of each potential response, not just how responses compare to one another in terms of preference. Towards this in our work, we create pairs consisting of an initial LLM-generated response and its refined counterparts that by nature are more contrastive (or closely comparable). Our results show that such a pairing results in improved alignment and agreement during preference ranking.
@@ 3. DESIGN CONSIDERATIONS TO IMPROVE AI WRITING
“The secret to good writing is good editing.It’s what separates hastily written, randomly punctuated, incoherent rants from learned polemics and op-eds, and cringe-worthy fan fiction from a critically acclaimed novel ” (Guinness, 2024). In this section, we outline the design principles and desiderata that guided our approach to improving AI writing through textual edits.
@@
This principle emphasizes creating a comprehensive taxonomy of edit categories (Faigley and Witte, 1981) rooted in an expert editor and writer’s practices. Prior work has shown that experts and novices define revising in very different ways with experts attending more systematically to different aspects of the text than novices (Bracewell et al., 1978; Sommers, 1980; Tyler et al., 1982; Rosson, 1983). By developing such a taxonomy, we aim to provide an approach to analyzing and enhancing LLM-generated text. It also allows for a more granular understanding of the specific areas where AI writing may fall short and enables targeted improvements. Sommers (1980) found that “experienced writers have a second objective; a concern for their readership”. Grounding the taxonomy in expert writing practices ensures that the edits align with the standards of high-quality writing and are acceptable to its readers. Finally, this principle also acknowledges the complexity of the editing process, recognizing that different categories of edits may be required at various levels of the text, from sentence-level corrections to broader structural changes (Hayes et al., 1987; Sommers, 1980; Yang et al., 2017).
@@
Preserving the core meaning and intent of the original text is crucial to maintaining coherence and faithfulness to the initial ideas. On the other hand, introducing substantive semantic changes is often required to adhere to the quality and characteristics of good writing. Prior work on edit taxonomies focuses on low-level syntactic operations (Faigley and Witte, 1981) or semantic edits (Yang et al., 2017; Daxenberger and Gurevych, 2013; Laban et al., 2023b) tailored to specific websites like Wikipedia. LLM-generated text often benefits from syntactic edits. These edits (primarily meaning preserving) enhance readability by diversifying sentence structures, expanding vocabulary choices, and minimizing repetitive phrasing. Consequently, semantic edits (both meaning preserving and changing) in AI writing are important for enhancing specificity or reducing unnecessary flourishes and clichés that can otherwise obscure meaning. Our methodology aims to navigate the tension between maintaining original meaning and introducing necessary improvements to mitigate AI-specific writing quirks.
@@
Current AI writing systems are developed using pre-trained language models (LMs) refined through human interaction, employing supervised learning and reinforcement learning (RL) techniques. Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019) is a key approach, transforming human input into training data to guide LMs toward desired outcomes. The most common type of feedback used with RLHF is binary preferences between pairs of examples sampled from one or more Language Models (Christiano et al., 2017). However, a learned preference ordering can fail to converge to the true one when the desirability of examples depends on noise (Gao et al., 2024a). Following recent work in preference learning (Dubey et al., 2024; D’Oosterlinck et al., 2024), we evaluate edits as a mechanism for enhancing alignment. An LLM-generated response that has been edited typically contains fewer undesirable traits and can be paired with the original LLM-generated response for preference ranking ($edited>original$).While contemporaneous works (Dubey et al., 2024; D’Oosterlinck et al., 2024) have conducted preliminary efforts to incorporate edits for improved preference data collection, we evaluate this approach in the context of creative writing.
@@ 4.1. Collecting Instruction and Response Pairs
Finally, we use the generated instructions to elicit responses from three state-of-the-art LLMs: OpenAI’s GPT-4o (OpenAI, 2024), Anthropic’s Claude-3.5-Sonnet (Anthropic, 2024), and Meta’s Llama 3.1-70b (Dubey et al., 2024). Each LLM is used to generate responses to one-third of our instruction data. We ensure that each LLM responds to instruction across all domains in equal proportion. To generate high-quality responses, we provide each LLM with the writing instruction, as well as the genre and source, and instruct it to adhere to the style of the venue. The prompt further specifies: “Try your best to be original, avoiding clichés or overused tropes. Do not use ornamental language and focus on nuance, simplicity, and subtext” (See Prompt in Appendix A Table 14). Through this process, we obtain 1,057 writing $<$instructions, response$>$ pairs, with responses averaging 205 words. This collection of instructions and LLM-generated responses serves as the foundation for the three studies we conducted: the formative study, the full-scale editing annotation, and the preference annotation.
@@ 4.4.6. Awkward Word Choice and Phrasing
Awkward phrasing can significantly reduce writing quality, often confusing or disengaging readers. This issue typically involves misused or disproportionate use of certain words (Kobak et al., 2024), unclear pronoun references, or an overuse of passive voice. In an email, W1 pointed out “Another little observation to share: a very common phrasing in these excerpts is ‘seem to _(verb)_’. This is not technically wrong it’s just inelegant, something many writing teachers have told me to avoid. Unless there is some specific uncertainty or doubt about the verb action, it’s always preferred to just use the verb without ‘seem’ (ex. from the current excerpt I have up: ‘amplified’ is better than ‘seemed to amplify’)”. Editing plays a crucial role in refining these elements. Through careful revision, writers can identify and replace imprecise or ill-fitting words with more appropriate alternatives, ensuring each term accurately conveys the intended meaning.